OpenStudy (anonymous):

if log base 7 of 3= 0.5645 and log base 7 of 8= 1.0686, what is the value of log base 7 of (3/8)?

OpenStudy (anonymous):

\[\log_{7}\frac{3}{8}=\log_{7}3-\log_{7}8\]

OpenStudy (anonymous):

substitute your values in there

OpenStudy (anonymous):

its called the quotient rule for logarithms

OpenStudy (anonymous):

let log base 7 of 3=log a and log base 7 of 8=log b log a/b = log a - log b then log base 7 of 3/8 = log base 7 of 3-log base 7 of 8

OpenStudy (anonymous):

thank you both